OpenStudy (anonymous):

Using complete sentences, explain which method you would use to solve the following system of equations and why you chose that method. Show your work and provide the solution to the system. 3x + 2y = 7 4x – 3y = –2

OpenStudy (anonymous):

substitiution!

OpenStudy (anonymous):

hold on i'll do it

OpenStudy (anonymous):

I actually used elimination. So first you have to get rid of a term, x or y "eliminate it" 3x + 2y = 7 4x – 3y = –2 im going to multiply to get both the x terms to cancel. Make one negative so they cancel. -4(3x + 2y = 7) 3(4x – 3y = –2) This Equals = -12x-8y= -28 12x-9y= -6 so you draw a line udner that and solve, you cross out the 12x's and solve for y. You get.... -17y= -34 y=2

OpenStudy (anonymous):

Now you take the y value which is 2 and plaug it back into one of the original equations. Where ever you see a y plugin 2. you try it

OpenStudy (anonymous):

Thank you.